# On Erlang mixture approximations for differential equations

### v2

@@ 5 Approximate system of ordinary differential equations
The block-diagonal elements, $A^{(i)}\in\mathbb{R}^{(M_{i}+1)\times(M_{i}+1)}$, $b^{(i)}\in\mathbb{R}_{\geq 0}^{M_{i}+1\times 1}$, and $c^{(i)}\in[0,1]^{1\times M_{i}+1}$, are given by
(75a) $A^{(i)}$ $=a_{i}(L^{(i)}-I^{(i)}),$ $b^{(i)}$ $=a_{i}e_{1}^{(i)},$ $c^{(i)}$ $=\begin{bmatrix}c_{0,i}&c_{1,i}&\cdots&c_{M_{i},i}\end{bmatrix}$, $i$ $=1,\ldots,n_{z},$
@@
where $I^{(i)},L^{(i)}\in\mathbb{R}^{M_{i}+1\times M_{i}+1}$ are an identity matrix and a lower shift matrix, respectively, i.e., $L^{(i)}$ contains ones in the first subdiagonal and zeros elsewhere. Furthermore, the first element of $e_{1}^{(i)}\in\mathbb{R}^{M_{i}+1\times 1}$ is one and the others are zero.
The transformation of the system of DDEs (71) to the system of ODEs (73) is standard [36, 41]. Therefore, we postpone the proof of Theorem 5.1 to Appendix A.
@@ 5.1 Steady states and stability criteria
In this section, we confirm that the steady state of the approximate system of ODEs satisfies the same equations as that of the original system, and we present the steady state stability criterion. In Section 5.2, we show that the roots involved in the criterion converge to those in the criterion for the original system of DDEs.
Theorem 5.2.
@@
(76) $0$ $=f(\bar{x},\bar{z}),$ $\bar{z}$ $=\bar{r}=h(\bar{x}),$
i.e., it satisfies the same equations as a steady state of the original system (2b)–(3).
Proof 5.3.
This result follows directly from the equivalence of the system of ODEs (73) to the system of DDEs (71) and Corollary 3.3. However, for completeness and to demonstrate some of the properties of the matrices $A$, $B$, and $C$, we prove it using direct computation. In steady state,
@@
(80) $\left(A^{(i)}\right)^{-1}$ $=\frac{-1}{a_{i}}\begin{bmatrix}1\\ 1&1\\ \vdots&&\ddots\\ 1&1&\cdots&1\end{bmatrix}$, $c^{(i)}\left(A^{(i)}\right)^{-1}b^{(i)}$ $=-\sum_{m=0}^{M_{i}}c_{m,i}=-1.$
Consequently, $CA^{-1}B=I$ is an identity matrix, and
(81) $\bar{z}$ $=\bar{r}=h(\bar{x}),$
which concludes the proof.
Theorem 5.4.
The system (73) is locally asymptotically stable around a steady state, $\bar{x}$, that satisfies (76) if $\operatorname{Re}\lambda<0$ for all eigenvalues $\lambda\in\mathbb{C}$ of the Jacobian matrix
(82) $J$ $=\begin{bmatrix}F&GC\\ BH&A\end{bmatrix}$,
where $F$, $G$, and $H$ are the Jacobian matrices given by (9). Furthermore, the eigenvalues, $\lambda$, that are not equal to $a_{i}$ for $i=1,\ldots,n_{z}$ will also satisfy the reduced characteristic equation
(83) $\hat{P}(\lambda)=\det(F-\lambda I+GQ(\lambda)H)$ $=0,$ $Q_{ij}(\lambda)$ $=\begin{cases}\sum\limits_{m=0}^{M_{i}}c_{m,i}\left(\frac{a_{i}}{a_{i}+\lambda}\right)^{m+1},&i=j,\\ 0,&i\neq j,\end{cases}$
@@
(84) $\det(J-\lambda I)$ $=\det\begin{bmatrix}F-\lambda I&GC\\ BH&A-\lambda I\end{bmatrix}$
$=\det\left(F-\lambda I-GC(A-\lambda I)^{-1}BH\right)\det(A-\lambda I),$
@@
Let each element of $\hat{\alpha}:\mathbb{R}\rightarrow\mathbb{R}^{n_{z}}$ be an Erlang mixture kernel. Then, the approximate characteristic function $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ and the auxiliary matrix $Q:\mathbb{C}\rightarrow\mathbb{C}^{n_{z}\times n_{z}}$ defined in (83) in Theorem 5.4 are equal to
(92) $\hat{P}(\lambda)$ $=\det(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\hat{\alpha}(s)\,\mathrm{d}sH),$ $Q_{ii}(\lambda)$ $=\int_{0}^{\infty}e^{-\lambda s}\hat{\alpha}_{i}(s)\,\mathrm{d}s,$
@@
Proof 5.7.
The result follows directly from the equivalence of the approximate systems of DDEs (71) and ODEs (73) and the stability criterion in Lemma 3.5. However, for completeness, we write out the proof in detail. First, we write out the following integral for an Erlang kernel of order $m$ with rate parameter $a_{i}$:
@@
(95) $G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\hat{\alpha}(s)\,\mathrm{d}sH$ $=G\operatorname{diag}\left\{\int_{0}^{\infty}e^{-\lambda s}\hat{\alpha}_{i}(s)\,\mathrm{d}s\right\}_{i=1}^{n_{z}}H=GQ(\lambda)H,$
@@
(96) $\hat{P}(\lambda)=\det(F-\lambda I+GQ(\lambda)H).$
Corollary 5.8.
If $\lambda=0$ is an eigenvalue of the Jacobian matrix $J$ in (82), it is also a solution to the characteristic equation (8).
Proof 5.9.
As the rate parameters, $a_{i}$, are positive for $i=1,\ldots,n_{z}$, an eigenvalue of zero must be a root of the reduced characteristic function in (83), i.e., it is not an eigenvalue of $A$. Furthermore,
(97) $Q_{ii}(0)$ $=\sum_{m=0}^{M_{i}}c_{m,i}=1,$
i.e., $Q(0)$ is an identity matrix. Consequently, for an eigenvalue of zero,
(98) $\det(F+GH)=0.$
Similarly, if $\lambda=0$, the characteristic equation (8) becomes
(99) $\det\left(F+G\int_{0}^{\infty}\operatorname{diag}\alpha(s)\,\mathrm{d}sH\right)$ $=\det(F+GH)=0,$
which, as we have just shown, is satisfied.
@@ 5.2 Convergence of the roots
In this section, we use Hurwitz’ convergence theorem [19, Thm. 5.6.4] to prove that the roots of the approximate characteristic function, $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ in Theorem 5.4, converge to those of $P:\mathbb{C}\rightarrow\mathbb{C}$ in Corollary 3.5 as the rate parameters in the Erlang mixture approximation goes to infinity. That is, we show that stability analyses based on the approximate system of ODEs in (73) will be accurate for a sufficiently large rate parameter.
@@
Let each element of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ be a regular and exponentially bounded kernel, i.e., let there exist $L,\rho\in\mathbb{R}_{>0}$ such that $\alpha_{i}(t)\leq Le^{-\rho t}$ for all $t\in\mathbb{R}_{\geq 0}$ and for $i=1,\ldots,n_{z}$. Furthermore, let the $i$’th element of $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ be an infinite-order Erlang mixture approximation with rate parameter $a_{i}\in\mathbb{R}_{>0}$ and coefficients defined by (17) based on $\alpha_{i}$. Then, for a given $\lambda_{\min}\in\mathbb{R}_{\leq 0}$ for which $\rho<\lambda_{\min}$, the approximate characteristic function $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ defined in (83) converges uniformly to the characteristic function $P:\mathbb{C}\rightarrow\mathbb{C}$ defined in (8) for every $\lambda\in\mathbb{C}$ for which $\operatorname{Re}\lambda\geq\lambda_{\min}$ as $a\rightarrow\infty$.
Proof 5.11.
@@
Here, we have used the expression for $Q_{ii}$ from Corollary 5.6. We consider the second integral first. The exponential boundedness of $\alpha$ implies that $\hat{\alpha}$ is also exponentially bounded (see Lemma 4.31) such that
@@
By assumption, $\lambda_{\min}>-\rho$, and according to Lemma 4.31, we can always choose $\bar{a}\in\mathbb{R}_{>0}$ such that $\bar{\rho}$ is arbitrarily close to $\rho$ for all $a_{i}>\bar{a}$, i.e., such that $\lambda_{\min}+\bar{\rho}>0$. Next, we bound the first integral, and we exploit the non-positivity of $\lambda_{\min}$:
@@
According to Theorem 4.37, we can choose $\bar{a}$ large enough that the last integral is arbitrarily small. Finally, we bound the norm difference between the characteristic matrices in the expressions for $\hat{P}$ and $P$ in (92) and (8), respectively, and we use the $\infty$-norm for convenience:
(103) $\left\|F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\hat{\alpha}(s)\,\mathrm{d}sH-\left(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\alpha(s)\,\mathrm{d}sH\right)\right\|_{\infty}$
$\leq\|G\|_{\infty}\int_{0}^{\infty}|e^{-\lambda s}|\|\operatorname{diag}(\hat{\alpha}(s)-\alpha(s))\|_{\infty}\,\mathrm{d}s\|H\|_{\infty}$
$=\|G\|_{\infty}\int_{0}^{\infty}e^{-\operatorname{Re}\lambda s}\max_{i=1,\ldots,n_{z}}|\hat{\alpha}_{i}(s)-\alpha_{i}(s)|\,\mathrm{d}s\|H\|_{\infty}.$
As we have shown above, for any $\lambda\in\mathbb{C}$ for which $\operatorname{Re}\lambda>\lambda_{\min}$, we can choose $s_{0}$ and $\bar{a}$ such that this expression is below any $\epsilon\in\mathbb{R}_{>0}$ whenever $a_{i}>\bar{a}$ for $i=1,\ldots,n_{z}$, which concludes the proof.
@@
Let the elements of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ be exponentially bounded and regular, i.e., let there exist $L,\rho\in\mathbb{R}_{>0}$ such that $\alpha_{i}(t)\leq Le^{-\rho t}$ for all $t\in\mathbb{R}_{\geq 0}$ and $i=1,\ldots,n_{z}$, and let the $i$’th element of $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ be an Erlang mixture kernel of infinite order with rate parameter $a_{i}\in\mathbb{R}_{>0}$ and coefficients defined by (17) based on $\alpha_{i}$. Then, the characteristic function $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ defined by (83) is holomorphic for $\lambda\in\mathbb{C}$ for which $\operatorname{Re}\lambda>-\rho$.
Proof 5.13.
@@
The derivative exists if the integral is finite. According to Lemma 4.31, $\hat{\alpha}_{i}$ is exponentially bounded if $\alpha_{i}$ is. Specifically, $\hat{\alpha}_{i}(t)\leq Le^{-\bar{\rho}t}$ for all $t$ where $0\leq\bar{\rho}<\rho$. Consequently,

### v3

@@ 4 Main theoretical results
The block-diagonal elements, $A^{(i)}\in\mathbb{R}^{(M_{i}+1)\times(M_{i}+1)}$, $b^{(i)}\in\mathbb{R}_{\geq 0}^{M_{i}+1\times 1}$, and $c^{(i)}\in\mathbb{R}^{1\times M_{i}+1}$, are given by
(22) $A^{(i)}$ $=a_{i}(L^{(i)}-I^{(i)}),$ $b^{(i)}$ $=a_{i}e_{1}^{(i)},$ $c^{(i)}$ $=\begin{bmatrix}c_{0,i}&c_{1,i}&\cdots&c_{M_{i},i}\end{bmatrix}$, $i$ $=1,\ldots,n_{z},$
@@
where $I^{(i)},L^{(i)}\in\mathbb{R}^{M_{i}+1\times M_{i}+1}$ are an identity matrix and a lower shift matrix, respectively, i.e., $L^{(i)}$ contains ones in the first subdiagonal and zeros elsewhere. Furthermore, the first element of $e_{1}^{(i)}\in\mathbb{R}^{M_{i}+1\times 1}$ is one and the others are zero, and $c_{m,i}\in\mathbb{R}$ is the $m$’th coefficient in the $i$’th Erlang mixture approximation. Then, the following asymptotic properties hold as $a_{i}\rightarrow\infty$ provided that $M_{i}/a_{i}\rightarrow\infty$ as well for $i=1,\ldots,n_{z}$.
  1. (a)
    The steady state of the ODEs (19) converges to that of the DDEs (3b)–(4) if the matrix $F+G\operatorname{diag}\beta(\infty)H$ is invertible, where $F$, $G$, and $H$ are given by (11).
  2. (b)
    For every isolated root, $\lambda\in\mathbb{C}$, of the characteristic function in (12), there are roots of the characteristic function for the ODEs (19) that converge to $\lambda$ provided that $\operatorname{Re}\lambda\geq\lambda_{\min}$ where $\rho_{i}<\lambda_{\min}\in\mathbb{R}_{\leq 0}$ for $i=1,\ldots,n_{z}$. The $i$’th rate parameter, $\rho_{i}\in\mathbb{R}_{>0}$, is obtained from the exponential bound on $\alpha_{i}$ (see Definition 3.3).
  3. (c)
    If the DDEs (3b)–(4) and ODEs (19) are locally asymptotically stable around their respective steady states, so is the dynamical system describing the state error, $x-\hat{x}$.
  4. (d)
    The solution, $\hat{x}$, to the IVP (19)–(20) involving ODEs, the memory state $\hat{z}$, and the delayed quantity $\hat{r}$ converge uniformly to the solution, $x$, to the IVP (3)–(4) involving DDEs, the memory state $z$, and the delayed quantity $r$, respectively, for $t\in(-\infty,t_{f}]$.
Corollary 4.5.
If the initial state function, $x_{0}$, is constant, then $\hat{z}_{mi}(t_{0})=\hat{r}_{i,0}$ where $\hat{r}_{0}=h(x_{0})$.
Proof 4.6.
The result follows from the fact that, by definition, the Erlang kernels in (13) integrate to one:
(23) $\hat{z}_{mi}(t_{0})$ $=\int_{-\infty}^{t_{0}}\ell_{mi}(t_{0}-s)\hat{r}_{i}(s)\,\mathrm{d}s=\hat{r}_{i,0}\int_{-\infty}^{t_{0}}\ell_{mi}(t_{0}-s)\,\mathrm{d}s=\hat{r}_{i,0}.$
In the remainder of this section, we present proofs of the statements in Proposition 4.2 and Theorem 4.4.
4.1 Regularity and convergence of Erlang mixture approximations
In this section, we prove that the Erlang mixture approximation defined in Proposition 4.2 is regular, that it converges uniformly and weakly (i.e., the integral converges), and that the integral of the absolute error converges to zero as the rate parameter, $a$, goes to infinity provided that the ratio $M/a$ also goes to infinity, i.e., that the order, $M$, approaches infinity faster than $a$.
Proof 4.7 (Proof of Proposition 4.2(a)).
The boundedness in Definition 3.2(a) follows directly from the substitution of the upper bound on $\alpha$ into the expression for the coefficients in (15):
(24) $|\hat{\alpha}(t)|$ $=\left|\sum_{m=0}^{M}\ell_{m}(t)\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s\right|\leq\sum_{m=0}^{M}\ell_{m}(t)\int_{s_{m}}^{s_{m+1}}K\,\mathrm{d}s=K\sum_{m=0}^{M}\ell_{m}(t)\Delta s$
$=K\sum_{m=0}^{M}\frac{(at)^{m}}{m!}ae^{-at}\frac{1}{a}=K\sum_{m=0}^{M}\frac{(at)^{m}}{m!}e^{-at}\leq Ke^{at}e^{-at}=K,$
where we have used the identity (2) and that all terms in the sum are positive, i.e., $\sum_{m=0}^{M}(at)^{m}/m!\leq\sum_{m=0}^{\infty}(at)^{m}/m!=e^{at}$ for all $a\in\mathbb{R}_{>0}$ and $M\in\mathbb{N}_{\geq 0}$. The continuity of $\hat{\alpha}$ in Definition 3.2(b) follows directly from the fact that it is a weighted sum of the Erlang kernels, $\ell_{m}$, which are continuous because they are products of monomials and exponential functions. Finally, we show that the integral of the absolute value of $\hat{\alpha}$ is finite as in Definition 3.2(c):
(25) $\int_{0}^{\infty}|\hat{\alpha}(s)|\,\mathrm{d}s$ $=\int_{0}^{\infty}\left|\sum_{m=0}^{M}c_{m}\ell_{m}(s)\right|\,\mathrm{d}s\leq\sum_{m=0}^{M}|c_{m}|\int_{0}^{\infty}\ell_{m}(s)\,\mathrm{d}s=\sum_{m=0}^{M}|c_{m}|$
$=\sum_{m=0}^{M}\left|\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s\right|\leq\int_{0}^{s_{M+1}}|\alpha(s)|\,\mathrm{d}s\leq\int_{0}^{\infty}|\alpha(s)|\,\mathrm{d}s=\rho(\infty)<\infty.$
Here, we have used Minkowski’s inequality and that the Erlang kernels are non-negative and integrate to one.
Proof 4.8 (Proof of Proposition 4.2(b)).
In order to prove that the Erlang mixture approximation $\hat{\alpha}$ converges uniformly to $\alpha$, we express it in terms of the delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, defined by (94) in Lemma B.1:
This equality holds according to Lemma B.1(a), and since $\delta_{a}$ satisfies Lemma B.1(b)(d), Lemma D.1 guarantees that $\hat{\alpha}$ converges uniformly to $\alpha$ for all $t\in\mathbb{R}_{\geq 0}$.
Proof 4.9 (Proof of Proposition 4.2(c)).
First, we bound the absolute difference between the two integrals:
(27) $\left|\int_{0}^{t}\hat{\alpha}(s)\,\mathrm{d}s-\int_{0}^{t}\alpha(s)\,\mathrm{d}s\right|$ $\leq\int_{0}^{t}|\hat{\alpha}(s)-\alpha(s)|\,\mathrm{d}s\leq t\sup_{s\in[0,t]}|\hat{\alpha}(s)-\alpha(s)|.$
For finite $t$, the rightmost expression goes to zero as $a\rightarrow\infty$ and $M/a\rightarrow\infty$. When $t$ is infinite, we split the integral at $t_{0}\in\mathbb{R}_{\geq 0}$:
(28) $\left|\int_{0}^{\infty}\hat{\alpha}(s)\,\mathrm{d}s-\int_{0}^{\infty}\alpha(s)\,\mathrm{d}s\right|$ $\leq\int_{0}^{t_{0}}|\hat{\alpha}(s)-\alpha(s)|\,\mathrm{d}s+\int_{t_{0}}^{\infty}|\hat{\alpha}(s)-\alpha(s)|\,\mathrm{d}s.$
As $\alpha$ is regular, the integral of its absolute value is finite. Furthermore, according to Lemma E.4, $\hat{\alpha}$ is tight, i.e., we can choose $t_{0}$ large enough that the second term is below $\epsilon/2$ for all values of $a$ and $M$. Next, we choose $a$ and $M$ large enough that the first term is below $\epsilon/2$ as well, which concludes the proof.
Proof 4.10 (Proof of Proposition 4.2(d)).
As $\alpha$ is regular, it is bounded. According to Proposition 4.2(a), this means that $\hat{\alpha}$ is also bounded because it is also regular. Consequently, according to Lemma E.2, $\hat{\alpha}$ is uniformly integrable. Furthermore, according to Lemma E.4 and Proposition 4.2(b), $\hat{\alpha}$ is also tight and converges pointwise to $\alpha$ (because it converges uniformly). Finally, according to Vitali’s convergence lemma [47, Sec. 19.1, p. 399], uniform integrability, tightness, and pointwise convergence implies that $\hat{\alpha}$ also converges in the $L^{1}$-norm, i.e., the stated result.
@@ 4.2 Steady states and stability criteria
In this section, we show that the steady state of the approximate system of ODEs converges to that of the original system, and we present the steady state stability criterion. In Section 4.3, we show that the roots involved in the criterion converge to those in the criterion for the original system of DDEs.
Lemma 4.11.
@@
(29) $0$ $=f(\bar{x},\bar{z}),$ $\bar{z}$ $=d\odot\bar{r},$ $\bar{r}$ $=h(\bar{x}),$
where $d\in\mathbb{R}^{n_{z}}$ and $d_{i}=\sum_{m=0}^{M_{i}}c_{m,i}$ for $i=1,\ldots,n_{z}$.
Proof 4.12.
In steady state,
@@
(33) $\left(A^{(i)}\right)^{-1}$ $=\frac{-1}{a_{i}}\begin{bmatrix}1\\ 1&1\\ \vdots&&\ddots\\ 1&1&\cdots&1\end{bmatrix}$, $c^{(i)}\left(A^{(i)}\right)^{-1}b^{(i)}$ $=-\sum_{m=0}^{M_{i}}c_{m,i}.$
Consequently, $CA^{-1}B=\operatorname{diag}d$, and
(34) $\bar{z}$ $=d\odot\bar{r},$ $\bar{r}$ $=h(\bar{x}).$
Proof 4.13 (Proof of Theorem 4.4(a)).
First, according to Proposition 4.2(c),
(35) $d_{i}=\sum_{m=0}^{M_{i}}c_{m,i}=\int_{0}^{\infty}\hat{\alpha}_{i}(s)\,\mathrm{d}s\rightarrow\beta_{i}(\infty)$
as $a_{i}\rightarrow\infty$ when $M_{i}/a_{i}\rightarrow\infty$ as well. Next, the implicit function theorem [41, Chap. 7, Thm. 3] guarantees that the solution to the steady state equations (8) for the DDEs (3b)–(4) is a locally unique and differentiable function if the Jacobian with respect to $\bar{x}$ is invertible. Consequently, in that case, the solution to the steady state equations (29) for the ODEs (19) converges to that of the DDEs (3b)–(4). Furthermore, the Jacobian is given by
(36) $\frac{\mathrm{d}{f}}{\mathrm{d}{\bar{x}}}(\bar{x},\bar{z})$ $=\frac{\partial{f}}{\partial{x}}(\bar{x},\bar{z})+\frac{\partial{f}}{\partial{z}}(\bar{x},\bar{z})\frac{\partial{\bar{z}}}{\partial{\bar{x}}},$ $\frac{\partial{\bar{z}}}{\partial{\bar{x}}}$ $=\operatorname{diag}\beta(\infty)\frac{\partial{\bar{r}}}{\partial{\bar{x}}},$ $\frac{\partial{\bar{r}}}{\partial{\bar{x}}}$ $=\frac{\partial{h}}{\partial{x}}(\bar{x}),$
and the matrix in Theorem 4.4(a) is obtained by substituting the matrices $F$, $G$, and $H$ from (11).
Lemma 4.14.
The system of approximate ODEs (19) is locally asymptotically stable around a steady state, $\bar{x}$, that satisfies (29) if $\operatorname{Re}\lambda<0$ for all eigenvalues $\lambda\in\mathbb{C}$ of the Jacobian matrix
(37) $J$ $=\begin{bmatrix}\hat{F}&\hat{G}C\\ B\hat{H}&A\end{bmatrix}$,
where $\hat{F}\in\mathbb{R}^{n_{x}\times n_{x}}$, $\hat{G}\in\mathbb{R}^{n_{x}\times n_{z}}$, and $\hat{H}\in\mathbb{R}^{n_{z}\times n_{x}}$ are the Jacobian matrices (11) evaluated in the steady state for the approximate system. Furthermore, the eigenvalues, $\lambda$, that are not equal to $a_{i}$ for $i=1,\ldots,n_{z}$ will also satisfy the reduced characteristic equation
[…]
@@
(39) $\det(J-\lambda I)$ $=\det\begin{bmatrix}\hat{F}-\lambda I&\hat{G}C\\ B\hat{H}&A-\lambda I\end{bmatrix}$
$=\det\left(\hat{F}-\lambda I-\hat{G}C(A-\lambda I)^{-1}B\hat{H}\right)\det(A-\lambda I),$
@@
Let each element of $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n_{z}}$ be an Erlang mixture kernel. Then, the approximate characteristic function $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ and the auxiliary matrix $Q:\mathbb{C}\rightarrow\mathbb{C}^{n_{z}\times n_{z}}$ defined in (38) in Lemma 4.14 are equal to
[…]
@@
Proof 4.17.
The result follows directly from the equivalence of the approximate systems of DDEs (152) and ODEs (19) and the stability criterion in Corollary 3.7. However, for completeness, we write out the proof in detail. First, we write out the following integral for an Erlang kernel of order $m$ with rate parameter $a_{i}$:
@@
(50) $\hat{G}\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\hat{\alpha}(s)\,\mathrm{d}s\hat{H}$ $=\hat{G}\operatorname{diag}\left\{\int_{0}^{\infty}e^{-\lambda s}\hat{\alpha}_{i}(s)\,\mathrm{d}s\right\}_{i=1}^{n_{z}}\hat{H}=\hat{G}Q(\lambda)\hat{H}$,
@@
[…]
@@ 4.3 Convergence of the roots
In this section, we use Hurwitz’ convergence theorem [19, Thm. 5.6.4] to prove that the roots of the approximate characteristic function, $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ in Lemma 4.14, converge to the isolated roots of $P:\mathbb{C}\rightarrow\mathbb{C}$ in Corollary 3.7 as the rate parameters and order in the Erlang mixture approximation go to infinity. That is, we show that stability analyses based on the approximate system of ODEs in (19) can be used to partially infer information about the stability of the original DDEs.
@@
Let each element of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n_{z}}$ be a regular and exponentially bounded kernel, i.e., let there exist $L,\rho\in\mathbb{R}_{>0}$ such that $|\alpha_{i}(t)|\leq Le^{-\rho t}$ for all $t\in\mathbb{R}_{\geq 0}$ and for $i=1,\ldots,n_{z}$. Furthermore, let the $i$’th element of $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n_{z}}$ be the Erlang mixture approximation of $\alpha_{i}$ of order $M_{i}\in\mathbb{N}_{\geq 0}$ with rate parameter $a_{i}\in\mathbb{R}_{>0}$. Then, for a given $\lambda_{\min}\in\mathbb{R}_{\leq 0}$ for which $\rho<\lambda_{\min}$, the approximate characteristic function $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ defined in (38) converges uniformly to the characteristic function $P:\mathbb{C}\rightarrow\mathbb{C}$ defined in (12) for every $\lambda\in\mathbb{C}$ for which $\operatorname{Re}\lambda\geq\lambda_{\min}$ as $a_{i}\rightarrow\infty$ if $M_{i}/a_{i}\rightarrow\infty$ as well for $i=1,\ldots,n_{z}$.
Proof 4.19.
@@
Here, we have used the expression for $Q_{ii}$ from Corollary 4.16. We consider the second integral first. The exponential boundedness of $\alpha$ implies that $\hat{\alpha}$ is also exponentially bounded (see Corollary 4.3) such that
@@
By assumption, $\lambda_{\min}>-\rho$, and according to Corollary 4.3, we can always choose $\bar{a}\in\mathbb{R}_{>0}$ and $\bar{M}\in\mathbb{N}_{\geq 0}$ (for which $\bar{M}/\bar{a}\rightarrow\infty$ as $\bar{a}\rightarrow\infty$) such that $\bar{\rho}$ is arbitrarily close to $\rho$ for all $a_{i}>\bar{a}$ and $M_{i}>\bar{M}$, i.e., such that $\lambda_{\min}+\bar{\rho}>0$. Next, we bound the first integral, and we exploit the non-positivity of $\lambda_{\min}$:
@@
According to Proposition 4.2(d) and because the integrand is non-negative, we can choose $\bar{a}$ and $\bar{M}$ (for which $\bar{M}/\bar{a}\rightarrow\infty$ as $\bar{a}\rightarrow\infty$) large enough that the last integral is arbitrarily small for $a_{i}>\bar{a}$ and $M_{i}>\bar{M}$. Finally, we bound the norm difference between the characteristic matrices in the expressions for $\hat{P}$ and $P$ in (47) and (12), respectively, and we use the $\infty$-norm for convenience:
(55) $\left\|\hat{F}-\lambda I+\hat{G}\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\hat{\alpha}(s)\,\mathrm{d}s\hat{H}-\left(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\alpha(s)\,\mathrm{d}sH\right)\right\|_{\infty}$
$=\Bigg\|\hat{F}-F+\hat{G}\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\hat{\alpha}(s)\,\mathrm{d}s(\hat{H}-H)$
$+\hat{G}\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}(\hat{\alpha}(s)-\alpha(s))\,\mathrm{d}sH+(\hat{G}-G)\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\alpha(s)\,\mathrm{d}sH\Bigg\|_{\infty}$
$\leq\|e_{F}\|_{\infty}+\|\hat{G}\|_{\infty}\int_{0}^{\infty}|e^{-\lambda s}|\|\operatorname{diag}\hat{\alpha}(s)\|_{\infty}\,\mathrm{d}s\|e_{H}\|_{\infty}$
$+\|\hat{G}\|_{\infty}\left\|\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}(\hat{\alpha}(s)-\alpha(s))\,\mathrm{d}s\right\|_{\infty}\|H\|_{\infty}+\|e_{G}\|\int_{0}^{\infty}|e^{-\lambda s}|\|\operatorname{diag}\alpha(s)\|_{\infty}\,\mathrm{d}s\|H\|_{\infty}$
Here, $e_{F}=\hat{F}-F\in\mathbb{R}^{n_{x}\times n_{x}}$, $e_{G}=\hat{G}-G\in\mathbb{R}^{n_{x}\times n_{z}}$, and $e_{H}=\hat{H}-H\in\mathbb{R}^{n_{z}\times n_{x}}$. As $f$, $g$, and $h$ are assumed to be continuously differentiable, the Jacobians, $F$, $G$, and $H$, are continuous and since the steady state of the approximate system converges to that of the original system as $a_{i}\rightarrow\infty$ and $M_{i}/a_{i}\rightarrow\infty$, the errors $e_{F}$, $e_{G}$, and $e_{H}$ converge to zero. The norm of the integral in the third term is
(56) $\left\|\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}(\hat{\alpha}(s)-\alpha(s))\,\mathrm{d}s\right\|_{\infty}$ $=\max_{i=1,\ldots,n_{z}}\left|\int_{0}^{\infty}e^{-\lambda s}(\hat{\alpha}_{i}(s)-\alpha_{i}(s))\,\mathrm{d}s\right|$
$\leq\max_{i=1,\ldots,n_{z}}\int_{0}^{\infty}|e^{-\lambda s}||\hat{\alpha}_{i}(s)-\alpha_{i}(s)|\,\mathrm{d}s.$
As we have shown above, for any $\lambda\in\mathbb{C}$ for which $\operatorname{Re}\lambda>\lambda_{\min}$, we can choose $s_{0}$, $\bar{a}$, and $\bar{M}$ (for which $\bar{M}/\bar{a}\rightarrow\infty$ as $\bar{a}\rightarrow\infty$) such that this expression is below any $\epsilon\in\mathbb{R}_{>0}$ whenever $a_{i}>\bar{a}$ and $M_{i}>\bar{M}$ for $i=1,\ldots,n_{z}$, which concludes the proof.
@@
Let the elements of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n_{z}}$ be exponentially bounded and regular, i.e., let there exist $L,\rho\in\mathbb{R}_{>0}$ such that $|\alpha_{i}(t)|\leq Le^{-\rho t}$ for all $t\in\mathbb{R}_{\geq 0}$ and $i=1,\ldots,n_{z}$, and let the $i$’th element of $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n_{z}}$ be the Erlang mixture approximation of $\alpha_{i}$ of order $M_{i}\in\mathbb{N}_{\geq 0}$ with rate parameter $a_{i}\in\mathbb{R}_{>0}$. Then, the characteristic function $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ defined by (38) is holomorphic for $\lambda\in\mathbb{C}$ for which $\operatorname{Re}\lambda>-\rho$.
Proof 4.21.
@@
The derivative exists if the integral is finite. According to Corollary 4.3, $\hat{\alpha}_{i}$ is exponentially bounded if $\alpha_{i}$ is. Specifically, $\hat{\alpha}_{i}(t)\leq Le^{-\bar{\rho}t}$ for all $t$ where $0\leq\bar{\rho}<\rho$. Consequently,